If a 1.00kHz sound source moves at a speed of 50.0m/s toward a listener who moves at a speed of 30.0m/s in a direction away from the source, what is the apparent frequency heard by the listener? The speed of sound in air is 343m/s.


1
Expert's answer
2021-02-19T18:59:24-0500

Applying the formula for the Doppler effect, obtain the observed frequency:


"f = \dfrac{c - v_l}{c - v_s}f_0"

where "c = 343 m\/s" is the speed of soung, "v_l = 30m\/s" is the speed of the listener, "v_s = 50m\/s" is the speed of the source, and "f_0 = 1kHz" is the natural frequency of the source. Thus, find


"f = \dfrac{343m/s - 30m/s}{343m/s - 50m/s}\cdot 1kHz \approx 1.068\space kHz"

Answer. 1.068 kHz.